OpenStudy (anonymous):

The cost in dollars to produce x cups of lemonade is represented by the function C(x) = 10 + 0.20x. The revenue in dollars earned by selling x cups of lemonade is represented by the function R(x) = 0.50x. What is the minimum number of whole cups of lemonade that must be sold for the revenue to exceed the cost?

OpenStudy (perl):

Please show any work you have done so far.

Vocaloid (vocaloid):

from the problem, we are given: revenue = R(x) = 0.50x cost = C(x) = 10 + 0.20x x = # of cups of lemonade for "revenue to exceed costs", this means revenue > costs, giving us the inequality 0.5x > 10 + 0.20x since we want to know the number of cups of lemonade that we need for revenue to exceed costs, we solve the inequality for x

OpenStudy (anonymous):

ok i got that also its just im stuck with the (>)

OpenStudy (anonymous):

Oh nevermind thank you!